# Split Unlearning

### v3

@@ 5.4 Label-related Privacy Analysis
Considering G3 (privacy-preserving label sharing), we evaluate the server’s ability to infer real label characteristics from the received masked labels and corresponding intermediate outputs. Given that the server lacks access to a local dataset and the shared labels are anonymized, it faces significant constraints in conducting semantic inference attacks on labels. We focus our analysis on the server’s capability to deduce the actual number of labels used by clients, which involves cracking the critical mapping mechanism $(\mathcal{G}_{Y},\mathcal{G}_{V}):\{Y_{q},V_{q}^{k}\}\rightarrow\{Y_{x}^{*},{V_{x}^{*}}^{k}\}$. This mapping is the key differentiator between SplitWiper and SplitWiper+, highlighting the enhanced privacy features of the former in protecting label information against server-based inference attacks.
In SplitWiper+, the server employs clustering algorithms on the received data pairs $\{Y_{x}^{*},{V_{x}^{*}}^{k}\}$ to group the labels with closer corresponding intermediate outputs into the same cluster, hypothesizing that these labels are expanded from the same original label. We assess the effectiveness of this clustering-based label matching attack using two metrics:
@@
Inclusive clustering accuracy assesses whether the distances among masked labels, derived from the same original label, are close enough to be accurately grouped by the clustering algorithm. This metric checks if related labels cluster together based on their proximity in feature space. In contrast, perfect clustering accuracy also considers whether there is sufficient separation between different groups, evaluating the algorithm’s ability to clearly distinguish between them. Together, these metrics offer insights into SplitWiper+’s effectiveness in preserving label-related privacy by maintaining the integrity of masked label structures during clustering.
@@
We employ the KMeans clustering algorithm [31], a standard method renowned for its robustness and widespread use in detecting patterns in ungrouped data. Given that the real number of labels, $Q$, is confidential to the server, we utilize the elbow method [31] to determine an appropriate number of clusters $K$. Additionally, to illustrate the potential privacy breach in SplitWiper+, we also conduct experiments where $Q$ is used directly as the parameter $K$ for KMeans. This scenario simulates the impact of a hypothetical leakage of $Q$, revealing how such exposure could compromise the privacy measures implemented within SplitWiper+.
@@
Given that the pre-trained models on each client are consistent across both VGG and ResNet18 architectures—since resource-limited clients have few layers and training techniques like dropouts and batch normalization are applied on the server side—the intermediate values show no significant variation. Therefore, we do not differentiate between VGG and ResNet18 in our experiments, focusing instead on the impact of intermediate values on label privacy. As shown in Fig. 5(a), when $Q$ is unknown and the server estimates the optimal cluster number using the elbow method, perfect clustering accuracy remains low, not exceeding 10% with $\gamma=3$. This suggests that SplitWiper+, using a label expansion strategy with $\gamma=3$, preserves over 90% of label-related privacy, effectively safeguarding sensitive information against potential inference attacks. As $\gamma$ increases, or with non-uniform expansion factors ($\gamma=\Gamma_{p}$), perfect clustering accuracy decreases, further enhancing privacy protection.
@@
In contrast, the inclusive clustering accuracy (Fig. 5(b)), is notably higher than the perfect clustering accuracy, due to its less stringent criteria. As more masked labels are clustered together by KMeans, the likelihood of satisfying the inclusive clustering accuracy increases. Therefore, under conditions of higher $\gamma$ or in SplitWiper with $\gamma=1$, when more masked labels are grouped into the same cluster, the inclusive clustering accuracy rises. However, this scenario primarily indicates that masked labels derived from the same original label are grouped together, without providing the means to differentiate between distinct groups of labels. While this may reveal minimal related information, it is insufficient for practical use.
@@
When the number of real labels $Q$ is disclosed to the server (Fig. 5(c)), clustering-based label matching attacks achieve higher perfect clustering accuracy, increasing by an average of 16% compared to when $Q$ remains confidential. This increase highlights the additional privacy risks posed by the exposure of label counts, underscoring the necessity of implementing privacy-preserving measures in SplitWiper+. Furthermore, a larger $\gamma$ exacerbates the risk related to the leakage of $Q$, as a higher expansion factor results in a greater number of masked labels derived from each original label, providing more information to assist clustering when $Q$ is known.
@@
Takeaways. SplitWiper+ protects label privacy and corresponding intermediate values through expansion and DP-based obfuscation, preventing server inference attacks via clustering. When the true label count is concealed, SplitWiper+ preserves over 90% of label privacy, and even when exposed, it still offers around 70% protection.
6 Related Work
Unlearning is challenging in distributed environments. Machine unlearning has not been fully explored within current distributed learning frameworks, including SL. While efficient unlearning strategies have gained attention in FL, which also involves collaboration between clients and a central server. Some researchers have worked on accelerating retraining across all clients, the most robust unlearning method in FL, using optimized algorithms like distributed Newton-type updates [32], SGA-EWC [33], and SFU [34]. Solutions like FedEraser [35] reduce the impact on non-involved clients through calibration but still require cooperation from some clients. However, in SL, where client resources are often limited, non-requesting clients may lack the incentive or ability to contribute, especially if they are uninterested or unable to participate due to network issues. This presents a significant challenge in SL, where developing unlearning methods that accommodate these constraints remains unresolved.
Existing SISA is inadequate to SL scenarios. SISA [6] is a state-of-the-art solution for robust and efficient machine unlearning, enabling rapid adaptation to data removal requests without requiring full model retraining. By strategically partitioning data into distinct “shards” linked to different model versions, SISA allows for the targeted unlearning of specific data shards without requiring the participation of all shards. This method has been successfully applied in various domains, including unlearning in graph neural networks [13] and large language models [36]. However, SISA-based unlearning faces significant challenges in multi-client SL scenarios due to the interdependence between client model segments and the shared server model segment, which violates the Isolated requirement of SISA. Consequently, the direct application of conventional SISA is impractical for existing SL scenarios.
Label privacy is important in SL unlearning. Beyond unlearning, privacy concerns regarding intermediate outputs and labels in SL have been widely studied. U-shaped SL [4] avoids label propagation to the server by placing the output layer on the client side but violates SISA’s Isolated requirement by necessitating the return of intermediate values across all clients, disabling the independent unlearning and adding overhead to resource-limited clients. If labels are shared but privacy is required, existing research primarily addresses privacy risks from transmitting intermediate outputs to the server. Wu et al.[37] introduced a DP mechanism to perturb these outputs, while Li et al.[20] developed the Marvell method to quantify privacy leakage and optimize noise structures. However, these methods overlook the exposure of sensitive information through label quantity and semantics. Xiao et al. [38] obfuscates both labels and outputs via multiple linear combinations, reducing association with original data but still fails to conceal label quantity, potentially exposing sensitive information.
Ours. Our design overcomes the mentioned challenges. SplitWiper adapts SISA for multi-client SL, starting with low-effort pre-training on the clients followed by freezing their weights. This allows for one-way-one-off propagation to the server, enabling efficient unlearning without involving non-participating clients. Additionally, SplitWiper+ enhances label privacy by expanding, shuffling, and anonymizing real labels, with intermediate outputs expanded and protected using a DP-based mechanism.
@@ 7 Conclusion
In this work, we achieve the first practical split unlearning framework. We presented SplitWiper and SplitWiper+, which designed an one-way-one-off propagation to decouple propagation between clients and between clients and the server, enabling SISA-based unlearning with absent clients. SplitWiper+ further enhanced client label privacy against the server by employing DP. Experiments across diverse data distributions and tasks demonstrated that SplitWiper effectively achieved unlearning by involving only the requesting clients, resulting in 0% unlearning accuracy and over 108% greater effectiveness on retained data compared to existing SL methods. Additionally, it maintained constant overhead and reduced computational and communication costs by over 99%. SplitWiper+ preserved over 90% of label privacy, effectively protecting sensitive information from the server.
@@ Appendix A Proof of Theorem 1
Appendix B Complexity Analysis
To further demonstrate the advantages of the SplitWiper and SplitWiper+ frameworks in reducing computational and communication overhead, we analyze its complexity in both learning and unlearning scenarios. We compare these metrics with traditional Vanilla and U-shaped SL, which adhere to the standard procedures of the usual round-robin SL processes detailed in [4], as shown in Table 1.
B.1 Learning Scenarios
B.1.1 Computational Complexity
Client-side computational complexity. In the SISA-based SplitWiper framework, the primary computational cost is derived from the pre-training of local models on each client-$k$, which is $O(N\cdot\mathcal{H}_{K})$, where $N$ is the epochs for client model pre-training and $\mathcal{H}_{K}$ is the complexity of a single training iteration. While in SplitWiper+, beyond the pre-training computational cost, there is an additional component that arises from the computations required by the label expansion strategy implemented after the pre-training phase on the client side. This additional cost is $O(\gamma\;\mathcal{H}_{exp})$, where $\mathcal{H}_{exp}$ represents the complexity of the label expansion-and-masking scheme, and $\gamma$ is the average of the label expansion factors.
Server-side computational complexity. In SplitWiper and SplitWiper+, the server’s updating of its model incurs a computational cost of $O(M\cdot\mathcal{H}_{S})$, where $M$ is the number of epochs for server model training and $\mathcal{H}_{S}$ is the complexity of a single server training iteration. This complexity mirrors that in the Vanilla SL framework, where the server-side complexity is also $O(M\cdot\mathcal{H}_{s})$. Conversely, the U-shaped SL framework generally exhibits slightly lower server-side complexity, which is $O(M\cdot(\mathcal{H}_{S}-\mathcal{H}_{out}))$, because clients undertake output layer operations, thus offloading some of the computational demand from the server.
B.1.2 Communication Overhead

### v4

@@ 6.4 Label-related Privacy Analysis
Considering G3 (privacy-preserving label sharing), we evaluate the server’s ability to infer real label characteristics, such as semantic information and label quantity, from the received masked labels and corresponding intermediate outputs. Given that the server lacks access to a local dataset and the shared labels are anonymized, it faces significant constraints in conducting semantic inference attacks on labels. We focus our analysis on the server’s capability to deduce the actual number of labels used by clients, as the protection of label quantity is a pivotal advancement in our approach, effectively preventing attackers from leveraging the exact number of labels to perform high-accuracy supervised learning attacks. The inference of the label quantity requires unraveling the critical mapping mechanism $(\mathcal{G}_{Y},\mathcal{G}_{V}):\{Y_{q},V_{q}^{k}\}\rightarrow\{Y_{x}^{*},{V_{x}^{*}}^{k}\}$–the key difference between SplitWiper+ and SplitWiper. This highlights the enhanced privacy features of the former in protecting label information against server-based inference attacks.
In SplitWiper+, the server, in its attempt to infer the actual label quantity, is limited to employing clustering algorithms on the received data pairs $\{Y_{x}^{*},{V_{x}^{*}}^{k}\}$. These algorithms group the labels with closer corresponding intermediate outputs into the same cluster, hypothesizing that these labels are expanded from the same original label. We assess the effectiveness of this clustering-based label matching attack using two metrics:
@@
Inclusive clustering accuracy evaluates whether masked labels originating from the same label are grouped correctly based on their proximity in feature space. In contrast, perfect clustering accuracy assesses both the separation between groups and the algorithm’s ability to distinguish them. These metrics provide insights into the effectiveness of SplitWiper+ in preserving label-related privacy while maintaining the structure of masked labels during clustering.
@@
We employ the KMeans clustering algorithm [53, 54], a standard method renowned for its robustness and widespread use in detecting patterns in ungrouped data. Given that the real number of labels, $Q$, is confidential to the server, we utilize the elbow method [55] to determine an appropriate number of clusters $K$. Additionally, to illustrate the potential privacy breach in SplitWiper+, we also conduct experiments where $Q$ is used directly as the parameter $K$ for KMeans. This scenario simulates the impact of a hypothetical leakage of $Q$, revealing how such exposure could compromise the privacy measures implemented within SplitWiper+ and further emphasizing the critical importance and value of concealing the explicit label quantity.
Given that the pre-trained models on each client are consistent across both VGG and ResNet18 architectures—owing to the limited number of layers on resource-constrained clients and the application of training techniques such as dropout and batch normalization on the server side—the intermediate values exhibit negligible variation. Consequently, we do not differentiate between VGG and ResNet18 in our experiments, focusing instead on the impact of intermediate values on label privacy.
@@
As illustrated in Fig.6(a), when $Q$ (the true number of clusters) is unknown and the server estimates the optimal number of clusters using the elbow method, perfect clustering accuracy remains low, not exceeding 10% for $\gamma=3$. This indicates that SplitWiper+, employing a label expansion strategy with $\gamma=3$, preserves over 90% of label-related privacy, effectively protecting sensitive information against inference attacks. Furthermore, as $\gamma$ increases or with non-uniform expansion factors ($\gamma=\Gamma_{p}$), perfect clustering accuracy decreases, providing enhanced privacy protection.
@@
In contrast, the inclusive clustering accuracy (Fig.6(b)), is notably higher than the perfect clustering accuracy, due to its less stringent criteria. As more masked labels are clustered together by KMeans, the likelihood of satisfying the inclusive clustering accuracy increases. Under conditions of higher $\gamma$ or with $\gamma=1$, when more masked labels are grouped into the same cluster, the inclusive clustering accuracy rises. However, this scenario primarily indicates that masked labels derived from the same original label are grouped together, without providing the means to differentiate between distinct groups of labels. While this may reveal minimal related information, it is insufficient for practical use.
@@
When the number of real labels $Q$ is disclosed to the server (Fig.6(c)), clustering-based label matching attacks achieve higher perfect clustering accuracy, increasing by an average of 16% compared to when $Q$ remains confidential. The increase highlights the additional privacy risks posed by the exposure of label counts, underscoring the necessity of implementing privacy-preserving measures in SplitWiper+. Furthermore, a larger $\gamma$ amplifies the risks associated with the leakage of $Q$. Higher expansion factors result in an increased number of masked labels derived from each original label, thereby providing additional information that facilitates clustering attacks when $Q$ is known.
@@
Takeaway (label privacy retained with one-way-one-off). SplitWiper+ protects label privacy and intermediate values through expansion and DP-based obfuscation, preventing server inference attacks via clustering techniques. It achieves over 90% privacy preservation when the true label count is concealed and approximately 70% even when the label count is exposed.
@@ 7 Conclusion
We achieved the first practical Split Unlearning framework with SplitWiper and SplitWiper+, which designed a one-way-one-off propagation to decouple propagation between clients and the server, enabling SISA unlearning with absent clients. SplitWiper+ further enhanced client label privacy against the server by employing DP. Experiments across diverse data distributions and tasks demonstrated that SplitWiper effectively achieved unlearning by involving only the requesting clients, resulting in 0% unlearning accuracy, and 8% better accuracy on retained data than non-SISA unlearning on existing SL frameworks. SplitWiper maintained constant overhead and reduced computational and communication costs by over 99%. SplitWiper+ preserved over 90% of label privacy, effectively protecting sensitive information from the server.
@@ Appendix A Proof of Theorem 1
Appendix B Complexity Analysis
To further demonstrate the advantages of the SplitWiper and SplitWiper+ frameworks in reducing computational and communication overhead, we analyze its complexity in both learning and unlearning scenarios. We compare these metrics with traditional Vanilla and U-shaped SL, which adhere to the standard procedures of the usual round-robin SL processes detailed in [4, 48], as shown in Table 5 and Table 6.
B.1 Learning Scenarios
B.1.1 Computational Complexity
Client-side computational complexity. In SplitWiper framework, the primary computational cost is derived from the pre-training of local models on each client $k$, which is $O(N\cdot\mathcal{H}_{K})$, where $N$ is the epochs for client model pre-training and $\mathcal{H}_{K}$ is the complexity of a single training iteration. While in SplitWiper+, beyond the pre-training computational cost, there is an additional component that arises from the computations required by the label expansion strategy implemented after the pre-training phase on the client side. This additional cost is $O(\gamma\;\mathcal{H}_{exp})$, where $\mathcal{H}_{exp}$ represents the complexity of the label expansion-and-masking scheme, and $\gamma$ is the average of the label expansion factors.
Table 5: Client-side Privacy and Overhead in SL
Framework Label Privacy Computational Communication
SplitWiper Low $O(\mathcal{H}_{K})$ $O(\mathcal{I}_{K}+\mathcal{I}_{Y})$
Table 6: Server-side Overhead in SL
Framework Computational Communication Storage
Learning Vanilla SL $O(M\cdot\mathcal{H}_{S})$ $O(M\cdot\mathcal{I}_{loss})$ $O(\mathcal{W}_{s})$
U-shaped SL $O(M\cdot(\mathcal{H}_{S}-\mathcal{H}_{out}))$ $O(M\cdot\mathcal{I}_{S})$ $O(\mathcal{W}_{s})$
SplitWiper $O(M\cdot\mathcal{H}_{S})$ NA $O(K\cdot\mathcal{D}_{inter}+\mathcal{W}_{s})$
SplitWiper+ $O(M\cdot\mathcal{H}_{S})$ NA $O(K\cdot\gamma\cdot\mathcal{D}_{inter}+\mathcal{W}_{s})$
Unlearning Vanilla SL $O(M\cdot K\cdot\mathcal{H}_{K})$ $O(M\cdot K\cdot(\mathcal{I}_{K}+\mathcal{I}_{Y}))$ $O(\mathcal{W}_{s})$
U-shaped SL $O(M\cdot K\cdot(\mathcal{H}_{K}+\mathcal{H}_{out}))$ $O(M\cdot K\cdot\mathcal{I}_{K})$ $O(\mathcal{W}_{s})$
SplitWiper $O(\mathcal{H}_{K})$ $O(\mathcal{I}_{K}+\mathcal{I}_{Y})$ $O(K\cdot\mathcal{D}_{inter}+\mathcal{W}_{s})$
SplitWiper+ $O(\mathcal{H}_{K}+\gamma\;\mathcal{H}_{exp})$ $O(\gamma\;(\mathcal{I}_{K}+\mathcal{I}_{Y}))$ $O(K\cdot\gamma\cdot\mathcal{D}_{inter}+\mathcal{W}_{s})$
  • $\mathcal{H}_{S}$: server-side training complexity; $\mathcal{H}_{out}$: output layer complexity; $\mathcal{I}_{loss}$: size of the server-side loss; $\mathcal{I}_{S}$: size of the server-side outputs; NA: no overhead applicable; $\mathcal{I}_{K}$: size of intermediate outputs; $\mathcal{I}_{Y}$: size of labels ; $\mathcal{W}_{s}$: size of server-side model weights; $\mathcal{D}_{inter}$: size of the client-side intermediate values.
Server-side computational complexity. The server’s computational cost for updating its model in SplitWiper and SplitWiper+ is $O(M\cdot\mathcal{H}{S})$, where $M$ denotes the number of training epochs, and $\mathcal{H}{S}$ represents the complexity of a single training iteration. This complexity mirrors that in the Vanilla SL framework, where the server-side complexity is also $O(M\cdot\mathcal{H}_{s})$. Conversely, the U-shaped SL framework generally exhibits slightly lower server-side complexity, which is $O(M\cdot(\mathcal{H}_{S}-\mathcal{H}_{out}))$, because clients undertake output layer operations, thus offloading some of the computational demand from the server.
B.1.2 Communication Overhead
Server-side communication complexity. In the proposed frameworks, the information transfer from clients to the server follows a one-way propagation manner, allowing us to consider that the server incurs no communication costs. In contrast, the Vanilla SL framework requires the server to send the loss back to each client to facilitate model training, leading to a communication complexity of $O(M\cdot\mathcal{I}_{loss})$ over $M$ epochs, where $\mathcal{I}_{loss}$ represents the size of the loss during each epoch. Similarly, our U-shaped SL necessitates the server to return outputs back to clients for operations related to the final output layer, resulting in a communication complexity of $O(M\cdot\mathcal{I}_{S})$, where $\mathcal{I}_{S}$ is the size of the server-side outputs.
B.1.3 Server-side Storage Overhead
In addition to computational and communication overheads, our SplitWiper and SplitWiper+ frameworks introduce additional storage overhead on the server to support the one-way-one-off propagation scheme. The server needs to store the intermediate values received from clients, ensuring that these values can be readily accessed during unlearning scenarios without disrupting other clients which do no request unlearning, whereas in Vanilla SL and U-shaped SL, storage overhead depends solely on $\mathcal{W}_{s}$, the size of the stored server-side model weights.
In SplitWiper, the storage complexity on the server is $O(K\cdot\mathcal{D}{inter}+\mathcal{W}_{s})$, where $K$ is the number of clients and $\mathcal{D}{inter}$ is the average size of the intermediate values transmitted by each client. In SplitWiper+, the intermediate values are expanded, resulting in a storage complexity of $O(K\cdot\gamma\cdot\mathcal{D}_{inter}+\mathcal{W}_{s})$ with the average of the label expansion factors $\gamma$.
B.2 Unlearning Scenarios
Conversely, SplitWiper and SplitWiper+ introduce an optimized approach, wherein unlearning requests are isolated to the specific client initiating the request, significantly reducing the overall burden. This isolation avoids the need for unrelated clients to engage in redundant computations and communications, thereby enhancing efficiency and reducing client-side total resource consumption. In terms of client-side computational complexity in SplitWiper framework, the primary costs come from the retraining of the local model on the initiating client $k$, which is $O(\mathcal{H}_{K})$. In contrast, in SplitWiper+, client $k$ also incurs the additional cost of performing a new round of the expansion-and-masking scheme, resulting in a computational complexity of $O(\mathcal{H}_{K}+\gamma\;\mathcal{H}_{exp})$.
The communication overhead is determined by transmitting the intermediate outputs from the initiating client $k$ to the server, leading to a complexity of $O(\gamma\;(\mathcal{I}_{K}+\mathcal{I}_{Y}))$ in the proposed frameworks, where $\gamma=1$ represents SplitWiper, and $\gamma>1$ represents SplitWiper+. Additionally, in unlearning scenarios, the server only updates the stored intermediate values related to client $k$ without requiring additional storage costs, allowing us to consider server-side storage overhead during the unlearning scenarios consistent with that of the learning scenarios. It satisfies G1 (independent unlearning), thus improving the robustness and reducing the overhead of unlearning processes.
@@ B.2 Unlearning Scenarios
Appendix C Algorithm of the Frameworks
The procedure steps in SplitWiper (cf. Sec.4) can be found from Algorithms 1&2, while that of its enhanced privacy-preserving option SplitWiper+ (cf. Sec.5) can be found from Algorithms 3&4.